Accelerator Update for April 8, 2009 - April 10, 2009 |
Wednesday April 8 The day shift began with the Tevatron (TeV) cooldown in progress and with F sector and the Transfer Hall being prepared for access, with Operations monitoring the Recycler stash (279.5E10) and the Antiproton Source (Pbar) stack (13.9E10), with the Linac, Booster, and Main Injector (MI) conducting maintenance, with the NuMI and MiniBooNE experiments conducting maintenance, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) conducting maintenance. Operations established beam to the Booster dump at 12:23 PM. Operations established beam to MiniBooNE at 12:26 PM. Operations established beam to NuMI and Pbar at 12:47 PM. Operations established beam to Switchyard at1:45 PM. Operations received permission to send beam to the Meson MTest T-953 experiment. Operations established beam to MTest at 3:26 PM and disabled their beam at 4 PM. Operations reported at 4:15 PM that the Recycler had stashed antiprotons once during the day shift. Operations put the TeV through a dry squeeze at 8:31 PM. At, 8:49 PM a TeV power supply (D0Q5) tripped off on a lead failure. Operators and Cryo System technicians changed the orifice. Operations put the TeV through a wet squeeze at 11:35 PM. Operations reported at 11:56 PM that the Recycler had stashed four times during the evening shift. Thursday April 9 The midnight shift began with the TeV in a wet squeeze, with Operations monitoring the stash (402.9E10), with Pbar stacking (97.4E10), and with NuMI and MiniBooNE taking beam. Operations put the TeV into shot setup at 1:03 AM. Operations established store 6961 at 2:17 AM with an initial luminosity of 255.97E30. Operations established beam to MTest at 6:09 AM. Operations reported at 7:56 AM that the Recycler had stashed four times during the midnight shift. Operations terminated store 6961 at 10:36 AM. Operations prepared the CDF and D0 collision halls for access at 10:52 AM. Operations put Pbar into standby for power supply work at 10:56 AM. CDF came out of access at 12:15 PM. D0 came out of access at 12:39 PM. Operations put the TeV through a dry squeeze at 1:10 PM. Operations put the TeV through a wet squeeze at 1:51 PM. Operations put the TeV through another wet squeeze at 2:48 PM. Operations put the TeV into shot setup at 3:10 PM. Operations prepared the Main Injector for access at 4:40 PM. At 4:52 PM, power was turned off to the MI-12 service building so that FESS could inspect the switch gear. Operations established store 6965 at 4:59 PM with an initial luminosity of 314.8E30. The Main Injector access began at 5:26 PM. The access is needed to fill the MI-30 Recycler Kicker with fluorinert. MI-12 power was restored at 6:18 PM. Main Injector access completed at 7:41 PM. Beam returned to NuMI and MiniBooNE at 7:45 PM. At 7:45 PM a ground fault indication was found on I:QF11B, power supply technicians investigated. PBar remained off during the repair. I:QF11B repair completed at 11:18 PM, PBar returned to stacking. Friday April 10 The midnight shift began with Operations monitoring the store (126.55E30) and stash (91.4E10), with Pbar stacking (16.9E10), and with NuMI and MiniBooNE taking beam. Operations established beam to MTest at 4:17 AM. The Plans for Friday and the Weekend The plans are to put a store in later this morning and stack, stash, and store throughout the weekend. |
|
More Information Comments and Suggestions |
| last modified 04/10/2009 email Fermilab |
| Security, Privacy, Legal | |